[Hallc_running] [New Logentry] RC update for Hall C (Sat-Sun)

mack at jlab.org mack at jlab.org
Sun Dec 1 18:00:02 EST 2019

Logentry Text:
 Work on the 3He target continues. The plan is for Hall C to be open Monday day for repairs, checklists, angle certs, and FSD updates. Then we lock up for hot checkout ~5pm, and after a few hours of hot checkout-ing, the MCC will drop the hall to Controlled Access for more target work. (Hence the target group gets a nice dinner at home, and will need to use keys when they return.)  

Shifts: There are virtual UNICORNS available for adoption later this month (ie, two blocks of DAY shifts). Very rare. Sign up at https://misportal.jlab.org/mis/apps/physics/shiftSchedule/index.cfm?experimentRunId=HALLC-POL3HE .

Accelerator: Making progress despite RF separator trips and Injector repeatedly dropping to power permit.  The operational restrictions for Hall C targets got updated. The procedure for ion chamber calibrations in Hall C will be updated (Spring 2019 targets are still assumed).

3He Target: Setup continues. Some control and cabling issues. The usual affirmations of our realism. Looks like at least a day-long to-do list  https://logbooks.jlab.org/entry/3745531 , https://logbooks.jlab.org/entry/3744714 .) JP will have to open the hall for the techs on Monday morning. This a CANS system configuration change which he controls. Most of us cannot now badge into the hall unless we're on a very short list. When it is safe to do so, JP will ensure the target is enclosed, etc,  and re-enable CANS for we the unwashed masses. 

Hall C infrastructure, lots to do: 

    Add FSDs for cooling jets and target motion. (Ethan B in conjunction with MCC) 

    Exercise the HMS magnet controls. Jack may have just resolved the waterflow interlocks on HMS Q2 and D power supplies. (me?)

    Resolve the possible lead flow issue on SHMS Q1. Just get us to 3 GeV/c for now.  (Mike F?) 

    Need to certify the SHMS angle down to 8.5deg, and the HMS down to 11.7deg. (Steve L and spotters?)

    The target guys need some support from Andy and Walter tracking down a leak, getting hydrogen, etc. https://logbooks.jlab.org/entry/3745256

    Go thru the alarm handler and resolve any issues. One outstanding item noted by Brad was the SHMS transfer line vacuum. https://logbooks.jlab.org/entry/3744714  (me?)

    Moller cool-down would be nice early this week (Dave G)

For the next few days, there will be an RC meeting every weekday at 4:15 pm on the 2nd floor of the Counting House. I probably won't spam you with a reminder. Bluejeans link: https://bluejeans.com/861439920 . 


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3745577
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20191201/a35dae17/attachment.html>

More information about the Hallc_running mailing list